OpenStudy (harrisonshark):

What is the range of the function y = x^2? all real numbers x ≥ 0 y ≥ 0

OpenStudy (harrisonshark):

please help

OpenStudy (anonymous):

domain of y=x^2 is negative infinity, infinity range is zero to infinity !

OpenStudy (anonymous):

Therefore, the range is: ALL REAL NUMBERS GREATER THAN OR EQUAL TO 0. or simply: 0 to +INFINITY.

OpenStudy (harrisonshark):

thx

OpenStudy (harrisonshark):

wait so B? or A?

OpenStudy (anonymous):

Here's the full explanation from the site (Read this and you got your answer!): "Hi there! the best way of solving this is picturing out what the graph might look like. Let's assume you had the graph of a parabola y=x^2. You know that for every x you substitute, there'd always be a value for y. Thus, the domain is ALL REAL NUMBERS or from -INFINITY to + INFINITY. The range on the other hand is different. We know that any number raised to the second power will always yield a positive integer or 0. Thus, y=x^2 won't have any negative y-values as the graph opens upward. Therefore, the range is: ALL REAL NUMBERS GREATER THAN OR EQUAL TO 0. or simply: 0 to +INFINITY. "

OpenStudy (harrisonshark):

got it now thank you

OpenStudy (anonymous):

Your welcome again!